If a car has a mass of 800kg and moves with a velocity if 25m/s what force is needed to stop the car in 50m
1
Expert's answer
2020-01-20T05:24:13-0500
"v^2=2as"

"F=ma=\frac{mv^2}{2s}"

"F=\frac{mv^2}{2s}=\frac{(800)(25)^2}{2(50)}=5000\ N"